# A closer look at some cyclic semifields

### v1

@@ Abstract.
We show that different choices of generators of the Galois group of $\mathbb{F}_{q^{n}}/\mathbb{F}_{q}$ ($n\geq 3$) produce non-isotopic cyclic semifields $\mathbb{F}_{q^{n}}[t;\sigma]/\mathbb{F}_{q^{n}}[t;\sigma](t^{m}-a)$ when $m=n$: for $n\geq m$, there are $\varphi(n)$ non-isomorphic classes of Sandler semifields $\mathbb{F}_{q^{n}}[t;\sigma]/\mathbb{F}_{q^{n}}[t;\sigma](t^{m}-a)$, one class for each generator $\sigma$ of ${\rm Gal}(\mathbb{F}_{q^{n}}/\mathbb{F}_{q})$ involved in their construction, where $\varphi$ is the Euler function. We prove that when $n=m$, two Sandler semifields constructed from different generators $\sigma_{1}$ and $\sigma_{2}$ of ${\rm Gal}(\mathbb{F}_{q^{n}}/\mathbb{F}_{q})$ are not isotopic. Hence when $n=m$ there are $\varphi(m)$ non-isotopic classes of these semifields, each class belonging to one choice of generator. We then present a full parametrization of the non-isomorphic Sandler semifields of order $q^{m^{2}}$ with center $\mathbb{F}_{q}$, and nuclei $\mathbb{F}_{q^{m}}$, when $m$ is prime and $\mathbb{F}_{q}$ contains a primitive $m$th root of unity. Since for $m=n$, two Sandler semifields constructed from the same generator are isotopic if and only if they are isomorphic, this parametrizes these Sandler semifields up to isotopy, and thus the corresponding non-Desarguesian projective planes as well. Most of our results are proved in all generality for any cyclic Galois field extension.
@@
Semifields, isotropy, isometry, cyclic semifields
@@ 1. Introduction
In this paper, we show that in certain cases, different choices of generators $\sigma_{1}$ and $\sigma_{2}$ of ${\rm Gal}(K/F)$ can produce non-isotopic Petit algebras. This is a consequence of our First Main Theorem 3.3:
@@
Let $n\geq 3$ and let $\sigma_{1}$ and $\sigma_{2}$ be any two distinct generators of the Galois group $\mathrm{Gal}(K/F)$. Suppose that $n\geq m$. Let $t^{m}-a_{1}\in K[t;\sigma_{1}]$ and $t^{m}-a_{2}\in K[t;\sigma_{2}]$, $a_{i}\in K\smallsetminus F$, both be not right-invariant. Then
@@
When $F=\mathbb{F}_{q}$ is a finite field, $n\geq m$, and both $t^{m}-a_{i}\in\mathbb{F}_{q^{n}}[t;\sigma_{i}]$, $i\in\{1,2\}$, are irreducible, then $\mathbb{F}_{q^{n}}[t;\sigma_{1}]/\mathbb{F}_{q^{n}}[t;\sigma_{1}](t^{m}-a_{1})$ and $\mathbb{F}_{q^{n}}[t;\sigma_{2}]/\mathbb{F}_{q^{n}}[t;\sigma_{2}](t^{m}-a_{2})$ are non-isomorphic semifields (Corollary 4.3). Semifields of the type $\mathbb{F}_{q^{n}}[t;\sigma]/\mathbb{F}_{q^{n}}[t;\sigma](t^{m}-a)$, $n\geq m$, are the opposite algebras (i.e., the dual semifields) of Sandler’s semifields [San1962]. In particular, hence there are $\varphi(m)$ ($\varphi$ the Euler function) different classes of non-isomorphic Sandler semifields that are the dual semifields of
@@
Hence for fixed $m\geq 3$, two proper semifields $(\mathbb{F}_{q^{m}}/\mathbb{F}_{q},\sigma_{1},a_{1})$ and $(\mathbb{F}_{q^{m}}/\mathbb{F}_{q},\sigma_{2},a_{2})$ are isotopic if and only if they are isomorphic.
When $m$ is prime and $\mathbb{F}_{q}$ contains a primitive root of unity, we count the proper nonassociative cyclic algebras (i.e., those Sandler semifields where $m=n$) up to isomorphism and thus up to isotropy, and give an explicit parametrization of the isomorphism (and thus, respectively, isotopism) classes of these algebras corresponding to a fixed $F=\mathbb{F}_{q}$. Since different choices of generators $\sigma_{1}$ and $\sigma_{2}$ of ${\rm Gal}(\mathbb{F}_{q^{m}}/\mathbb{F}_{q})$ produce non-isomorphic nonassociative cyclic algebras of degree $m$, there are more non-isomorphic Sandler semifields than previously counted.
The paper is organized as follows. After the preliminaries in Section 2, Section 3 contains the work towards our First Main Theorem, cited above. In Section 4, we prove our Second Main Theorem 4.1, count the number of Sandler semifields that are opposite algebras of nonassociative cyclic algebras of prime degree $m$ (Theorem 4.6) and give a parametrization of the algebras $(\mathbb{F}_{q^{m}}/\mathbb{F}_{q},\sigma_{1},a_{1})$ (Theorems 5.1 and 5.2) under the assumption that the base field $F$ contains a primitive $m$th root of unity in Section 5. Since two proper semifields coordinatize the same non-Desarguesian projective plane if and only if they are isotopic, we have thus also obtained a parametrization of the non-Desarguesian projective planes related to $(\mathbb{F}_{q^{m}}/\mathbb{F}_{q},\sigma_{1},a_{1})$.
We finish with a brief outlook on some open questions and what our results mean for linear codes in Section 6.
@@ 2.1. Nonassociative algebras
An algebra $A\not=0$ is called a division algebra if for any $a\in A$, $a\not=0$, the left multiplication with $a$, $L_{a}(x)=ax$, and the right multiplication with $a$, $R_{a}(x)=xa$, are bijective. If it is not clear which algebra structure on $V$ is considered, we denote the left and right multiplication with $a\in V$ by $L_{A}(a)$ and $R_{A}(a)$. If $A$ has finite dimension over $F$, $A$ is a division algebra if and only if $A$ has no zero divisors [Sch1995, pp. 15, 16].
@@ 2.2. Twisted polynomial rings
The twisted polynomial ring $R=K[t;\sigma]$ is the ring of polynomials $\{a_{0}+a_{1}t+\dots+a_{m}t^{m}\,|\,a_{i}\in K\}$, with term-wise addition and multiplication given by the rule $ta=\sigma(a)t$ for all $a\in K$ [Ore1933]; see also [Jac1996, Ch I]. The constant nonzero polynomials $K^{\times}$ are the units of $R$.
@@ 2.3. Petit algebras, nonassociative cyclic algebras and Sandler’s semifields
In this paper we will investigate the case when $f(t)=t^{m}-a\in K[t;\sigma]$ is not right-invariant, which is the case if and only if $a\in K\smallsetminus F$ [Pe1966]. The algebra $K[t;\sigma]/K[t;\sigma](t^{m}-a)$ is a division algebra over $F$ if and only if $f(t)=t^{m}-a\in K[t;\sigma]$ is irreducible [Pe1966, (7)]. We also know that $A=K[t;\sigma]/K[t;\sigma](t^{m}-a)$ is a division algebra if and only if ${\rm Nuc}_{r}(A)$ is a division algebra, e.g. see [GoTor2014, Proposition 4]. Over finite base fields and for $m\leq n$, the opposite algebras of $K[t;\sigma]/K[t;\sigma](t^{m}-a)$ are called Sandler’s semifields.
@@
Let us turn to the question of when two such algebras are isomorphic.
@@
Let $K/F$ be a cyclic Galois extension of $F$ of degree $n$ and let $\sigma$ be a generator of $\mathrm{Gal}(K/F)$. Let $a,b\in K\smallsetminus F$. For $n\geq m$, we have
@@
For $n<m$, in Equation 2.1 holds for some $\tau\in\mathrm{Gal}(K/F)$ and some $k\in K^{\times}$, then
@@
Let $K/F$ be a cyclic Galois extension of $F$ of degree $n\geq m$ and let $\sigma$ be a generator of $\mathrm{Gal}(K/F)$. If $a,b\in K\smallsetminus F$ such that
@@
[Br2018, Theorem 3.11] Suppose either that $m=2$ or $3$, or that $m\geq 5$ is a prime such that $F$ contains a primitive $m$th root of unity. Let $a\in K\smallsetminus F$. Then $K[t;\sigma]/K[t;\sigma](t^{m}-a)$ is a division algebra if and only if
@@
Suppose either that $m=2$ or $3$, or that $m\geq 5$ is a prime such that $F$ contains a primitive $m$th root of unity.
@@
(ii) If $F$ is a finite field and $N_{K/F}(a)\not\in(F^{\times})^{m}$, then $K[t;\sigma]/K[t;\sigma]f$ is a division algebra.
@@
We note that $n$ need not be equal to $m$ here.
For $m=n$, the unital central nonassociative algebra $K[t;\sigma]/K[t;\sigma](t^{m}-a)$ is called a nonassociative cyclic algebra of degree $n$ and denoted by $(K/F,\sigma,a)$ (cf. [Ste2014] who studies the opposite algebras, however). When $a\in F^{\times}$, the associative algebra $(K/F,\sigma,a)$ is a classical associative central simple cyclic algebra of degree $m$ [Jac1996, p. 19]. For $n=m$ we also have a stronger result than Corollary 2.6:
@@ 3. Different generators of $\mathrm{Gal}(K/F)$ yield nonisomorphic Sandler algebras
Let $F$ be an arbitrary field such that $F$ admits a cyclic Galois extension $K$ of degree $n>2$ and let $\sigma$ be a generator of the Galois group of $K/F$. We will now generalize Theorem [NevPu2024, Theorem 5.1] which covers the case $n=m$ that occurs when studying nonassociative cyclic algebras $(K/F,\sigma,a)$ to the case that $n\geq m$, starting with some general observations.
@@
Let $\sigma_{1}$ and $\sigma_{2}$ be any two distinct generators of $\mathrm{Gal}(K/F)$. There exists $1<j<n$, with $(j,n)=1$, such that $\sigma_{1}=\sigma_{2}^{j}$. Let $f\in K[t;\sigma_{1}]$ and $g\in K[t;\sigma_{2}]$ be both not right-invariant of degree $m$, and suppose that $n\geq m$. Let
@@
for $i\in\{0,\ldots,m-1\}$. Since $\sigma_{1}=\sigma_{2}^{j}$ and $\sigma_{i}$ has order $n\geq m$ it follows that $H(t)=kt^{j}$ for some $k\in K$. Since $H$ is an isomorphism, we know $k\in K^{\times}$. ∎
@@
Let $\sigma_{1}$ and $\sigma_{2}$ be any two distinct generators of $\mathrm{Gal}(K/F)$. Let $t^{m}-a_{1}\in K[t;\sigma_{1}]$ and $t^{m}-a_{2}\in K[t;\sigma_{2}]$ be not right-invariant, and suppose that $n\geq m$. Then
@@
Since $\sigma_{1},\sigma_{2}$ both generate $\mathrm{Gal}(K/F)$, there is some $1<j<n$, with $(j,n)=1$, such that $\sigma_{1}=\sigma_{2}^{j}$. Let $\ell$ be the least positive integer such that $\ell j>m$; since $\sigma_{1}\neq\sigma_{2}$, $\ell<m$. Suppose that $H:K[t;\sigma_{1}]/K[t;\sigma_{1}](t^{m}-a_{1})\rightarrow K[t;\sigma_{2}]/K[t;\sigma_{2}](t^{m}-a_{2})$ is an $F$-algebra isomorphism. Then $H|_{K}=\tau\in{\rm Gal}(K/F)$ and $H(t)=kt^{j}$ for some $k\in K^{\times}$ by Lemma 3.2. Since the Galois group is cyclic, and thus abelian, $\tau$ commutes with $\sigma_{i}$ on $K$, for $i\in\{1,2\}$. Let us rename the indeterminants in both algebras to $t_{1}$ and $t_{2}$, respectively, for a clearer argument. For any $s\geq 1$, put $\gamma_{s}(k)=k\sigma_{2}^{j}(k)\sigma_{2}^{2j}(k)\cdots\sigma_{2}^{(s-1)j}(k)$. Then one can show by induction that
$H(t_{1}^{s})=\begin{cases}\gamma_{s}(k)t_{2}^{sj}&\text{if $1\leq s<\ell$;}\\ \gamma_{\ell}(k)\sigma^{\ell j-m}(a_{2})t_{2}^{\ell j-m}&\text{if $s=\ell$.}\end{cases}$
@@
$t_{1}\cdot t_{1}^{\ell}=t_{1}^{\ell}\cdot t_{1}=\begin{cases}t_{1}^{\ell+1}&\text{if $\ell+1<m$};\\ a_{1}&\text{if $\ell+1=m$.}\end{cases}$
@@
$H(t_{1}^{\ell})H(t_{1})=\gamma_{\ell}(k)\sigma^{\ell j-m}(a_{2})t_{2}^{\ell j-m}\cdot kt_{2}^{j}$
@@
$H(t_{1}^{\ell})H(t_{1})=\begin{cases}\gamma_{\ell+1}(k)\sigma^{\ell j-m}(a_{2})t_{2}^{\ell j+j-m}&\text{if $(\ell+1)j<2m$;}\\ \gamma_{\ell+1}(k)\sigma^{\ell j-m}(a_{2})\sigma^{{\ell j+j-2m}}(a_{2})t_{2}^{\ell j+j-2m}&\text{if $2m\leq(\ell+1)j$}\end{cases}$
@@
$H(t_{1})H(t_{1}^{\ell})=kt_{2}^{j}\cdot\gamma_{\ell}(k)\sigma_{2}^{\ell j-m}(a_{2})t_{2}^{\ell j-m},$
@@
$H(t_{1})H(t_{1}^{\ell})=\begin{cases}\gamma_{\ell+1}(k)\sigma_{2}^{\ell j-m+j}(a_{2})t_{2}^{\ell j+j-m}&\text{if $(\ell+1)j<2m$;}\\ \gamma_{\ell+1}(k)\sigma_{2}^{\ell j-m+j}(a_{2})\sigma_{2}^{\ell j-2m+j}(a_{2})t_{2}^{\ell j+j-2m}&\text{if $2m\leq(\ell+1)j$}.\end{cases}$
@@
Let $n\geq m$. Let $\varphi$ be the Euler function. Then there are $\varphi(n)$ different classes of proper Petit algebras of type $K[t;\sigma_{i}]/K[t;\sigma_{i}](t^{m}-a)$; one class for each choice of generator $\sigma_{i}$ of $\mathrm{Gal}(K/F)$. Algebras in two different classes are not isomorphic.
The assumption that $n\geq m$ in Lemma 3.2 is indeed needed in the proof for (3.3). If $n<m$ we obtain a weaker version:
@@
Let $F$ be an arbitrary field and let $n\geq 3$ be the degree of a cyclic Galois extension $K/F$. Let $\sigma_{1}$ and $\sigma_{2}$ be any two distinct generators of the Galois group $\mathrm{Gal}(K/F)$. Since $\sigma_{1},\sigma_{2}$ both generate $\mathrm{Gal}(K/F)$, there is some $1<j<n$, with $(j,n)=1$, such that $\sigma_{1}=\sigma_{2}^{j}$. Let $f\in K[t;\sigma_{1}]$ and $g\in K[t;\sigma_{2}]$ be both not right-invariant of degree $m$, $n<m$. Let

### v2

@@ Abstract.
We show that different choices of generators $\sigma$ of the Galois group of $\mathbb{F}_{q^{n}}/\mathbb{F}_{q}$ produce non-isomorphic cyclic semifields $\mathbb{F}_{q^{n}}[t;\sigma]/\mathbb{F}_{q^{n}}[t;\sigma](t^{m}-a)$ when $n\geq m-1$: there are thus $\varphi(n)$ non-isomorphic classes of Sandler semifields $\mathbb{F}_{q^{n}}[t;\sigma]/\mathbb{F}_{q^{n}}[t;\sigma](t^{m}-a)$, one class for each generator $\sigma$ involved in their construction, where $\varphi$ is the Euler function. We prove that when $n=m$, two Sandler semifields constructed from different generators $\sigma_{1}$ and $\sigma_{2}$ of ${\rm Gal}(\mathbb{F}_{q^{n}}/\mathbb{F}_{q})$ are not isotopic. Hence when $n=m$ there are $\varphi(m)$ non-isotopic classes of these semifields, each class belonging to one choice of generator. We then present a full parametrization of the non-isomorphic Sandler semifields $\mathbb{F}_{q^{m}}[t;\sigma]/\mathbb{F}_{q^{m}}[t;\sigma](t^{m}-a)$ , when $m$ is prime and $\mathbb{F}_{q}$ contains a primitive $m$th root of unity. Since for $m=n$, two Sandler semifields constructed from the same generator are isotopic if and only if they are isomorphic, this parametrizes these Sandler semifields up to isotopy, and thus parametrizes both the corresponding non-Desarguesian projective planes, and maximum rank distance codes. Most of our results are proved in all generality for any cyclic Galois field extension.
@@
Semifields, isotropy, isometry, cyclic semifields, MRD codes
@@ 1. Introduction
In this paper, we show that in certain cases, different choices of generators $\sigma_{1}$ and $\sigma_{2}$ of ${\rm Gal}(K/F)$ indeed produce non-isotopic Petit algebras. This is a consequence of our First Main Theorem 3.3:
@@
Let $n\geq 3$ and let $\sigma_{1}$ and $\sigma_{2}$ be any two distinct generators of the Galois group $\mathrm{Gal}(K/F)$. Suppose that $n\geq m-1$. Let $t^{m}-a_{1}\in K[t;\sigma_{1}]$ and $t^{m}-a_{2}\in K[t;\sigma_{2}]$, $a_{i}\in K\smallsetminus F$, both be not right-invariant. Then
@@
When $F=\mathbb{F}_{q}$ is a finite field, $n\geq m-1$, and both $t^{m}-a_{i}\in\mathbb{F}_{q^{n}}[t;\sigma_{i}]$, $i\in\{1,2\}$, are irreducible, then $\mathbb{F}_{q^{n}}[t;\sigma_{1}]/\mathbb{F}_{q^{n}}[t;\sigma_{1}](t^{m}-a_{1})$ and $\mathbb{F}_{q^{n}}[t;\sigma_{2}]/\mathbb{F}_{q^{n}}[t;\sigma_{2}](t^{m}-a_{2})$ are non-isomorphic semifields (Corollary 4.3). Semifields of the type $\mathbb{F}_{q^{n}}[t;\sigma]/\mathbb{F}_{q^{n}}[t;\sigma](t^{m}-a)$, $n\geq m-1$, are the opposite algebras (i.e., the dual semifields) of Sandler’s semifields [San1962]. In particular, hence there are $\varphi(m)$ ($\varphi$ the Euler function) different classes of non-isomorphic Sandler semifields that are the dual semifields of
@@
Hence for fixed $m\geq 3$, in particular two proper semifields $(\mathbb{F}_{q^{m}}/\mathbb{F}_{q},\sigma_{1},a_{1})$ and $(\mathbb{F}_{q^{m}}/\mathbb{F}_{q},\sigma_{2},a_{2})$ are isotopic if and only if they are isomorphic.
When $m=2$ or when $m>2$ is prime and $\mathbb{F}_{q}$ contains a primitive root of unity, we count the proper nonassociative cyclic algebras (i.e., those Sandler semifields where $m=n$) up to isomorphism and thus up to isotropy, and give an explicit parametrization of the isomorphism (and thus, respectively, isotopism) classes of these algebras corresponding to a fixed $F=\mathbb{F}_{q}$ (Theorems 5.1 and LABEL:). Since different choices of generators $\sigma_{1}$ and $\sigma_{2}$ of ${\rm Gal}(\mathbb{F}_{q^{m}}/\mathbb{F}_{q})$ produce non-isomorphic nonassociative cyclic algebras of degree $m$, there are more non-isomorphic Sandler semifields than previously counted. Since two proper semifields coordinatize the same non-Desarguesian projective plane if and only if they are isotopic, we have thus also obtained a parametrization of the non-Desarguesian projective planes related to $(\mathbb{F}_{q^{m}}/\mathbb{F}_{q},\sigma_{1},a_{1})$ up to isomorphism.
The paper is organized as follows. After the preliminaries in Section 2, Section 3 contains the work towards our First Main Theorem, cited above. In Section 4, we prove our Second Main Theorem 4.1, count the number of Sandler semifields that are opposite algebras of nonassociative cyclic algebras of prime degree $m$ (Theorem 4.5) and give a parametrization of the algebras $(\mathbb{F}_{q^{m}}/\mathbb{F}_{q},\sigma_{1},a_{1})$ (Theorems 5.1 and 5.2) under the assumption that for $m>2$, the base field $F$ contains a primitive $m$th root of unity in Section 5.
We finish with a brief outlook on some open questions and what our results mean for linear codes in Section 6:
We prove that the maximum rank distance (MRD) codes obtained from proper nonassociative Petit division algebras of the kind $\mathbb{F}_{q^{m}}[t,\sigma]/\mathbb{F}_{q^{m}}[t,\sigma](t^{m}-a)$ depend on the choice of the generator $\sigma$ of the Galois group: different choices of generators yield a total of $\varphi(m)$ different classes of non-equivalent $\mathbb{F}_{q}$-linear MRD-codes in $M_{m^{2}}(\mathbb{F}_{q})$, respectively in $M_{m}(\mathbb{F}_{q^{m}})$ (Theorems 6.2 and 6.3).
In all of the existing literature, the question whether such a code depends on the choice of its generator $\sigma$ was never considered. This result shows how important this choice actually is.
Moreover, we now have a full parametrization of these codes when $m=2$ or when $m>2$ and $\mathbb{F}_{q}$ contains a primitive $m$th root of unity $\zeta$: each generator $\sigma$ and each $a\in\mathcal{S}_{2}(\mathbb{F}_{q^{2}})$, respectively each $a\in\mathcal{S}(\mathbb{F}_{q^{m}})$, yields a different (as in non-equivalent) MRD-code in $M_{m^{2}}(\mathbb{F}_{q})$, respectively in $M_{m}(\mathbb{F}_{q^{m}})$ (Theorem 6.3).
We conclude with an application of our First Main Theorem to skew $\sigma$-constacyclic codes: Let $n\geq 3$, $n\geq m-1$, and let $\sigma_{1}$ and $\sigma_{2}$ be any two distinct generators of ${\rm Gal}(\mathbb{F}_{q^{n}}/\mathbb{F}_{q})$. Let $a_{i}\in K\smallsetminus F$, then $a_{1}$ and $a_{2}$ are not $(m,\sigma)$-equivalent and not $(m,\sigma)$-isometric (Theorem 6.4).
@@ 2.1. Nonassociative algebras
An algebra $A\not=0$ is called a division algebra if for any $a\in A$, $a\not=0$, the left multiplication with $a$, $L_{a}(x)=ax$, and the right multiplication with $a$, $R_{a}(x)=xa$, are bijective. If $A$ has finite dimension over $F$, $A$ is a division algebra if and only if $A$ has no zero divisors [Sch1995, pp. 15, 16].
@@ 2.2. Twisted polynomial rings
The twisted polynomial ring $R=K[t;\sigma]$ is the ring of polynomials $\{a_{0}+a_{1}t+\dots+a_{m}t^{m}\,|\,a_{i}\in K\}$, with term-wise addition and multiplication given by the rule $ta=\sigma(a)t$ for all $a\in K$ [Ore1933]; see also [Jac1996, Chapter I]. The constant nonzero polynomials $K^{\times}$ are the units of $R$.
@@ 2.3. Petit algebras, nonassociative cyclic algebras and Sandler’s semifields
In this paper we will investigate the case when $f(t)=t^{m}-a\in K[t;\sigma]$ is not right-invariant, which is the case if and only if $a\in K\smallsetminus F$ [Pe1966]. The algebra $K[t;\sigma]/K[t;\sigma](t^{m}-a)$ is a division algebra over $F$ if and only if $f(t)=t^{m}-a\in K[t;\sigma]$ is irreducible [Pe1966, (7)]. We also know that $A=K[t;\sigma]/K[t;\sigma](t^{m}-a)$ is a division algebra if and only if ${\rm Nuc}_{r}(A)$ is a division algebra, e.g. see [GoTor2014, Proposition 4]. Over finite base fields and for $n\geq m$, the opposite algebras of $K[t;\sigma]/K[t;\sigma](t^{m}-a)$ are called Sandler’s semifields.
@@
Let us turn to the question of when two such algebras are isomorphic:
@@
Let $K/F$ be a cyclic Galois extension of $F$ of degree $n$ and let $\sigma$ be a generator of $\mathrm{Gal}(K/F)$. Let $a,b\in K\smallsetminus F$. For $n\geq m-1$, we have
@@
For $n<m-1$, if Equation 2.1 holds for some $\tau\in\mathrm{Gal}(K/F)$ and some $k\in K^{\times}$, then
@@
Let $K/F$ be a cyclic Galois extension of $F$ of degree $n\geq m-1$ and let $\sigma$ be a generator of $\mathrm{Gal}(K/F)$. If $a,b\in K\smallsetminus F$ such that
@@
[Br2018, Theorem 3.11] Let $K/F$ be a cyclic Galois extension of $F$ of degree $n$ and let $\sigma$ be a generator of $\mathrm{Gal}(K/F)$. Suppose either that $m=2$ or $3$, or that $m\geq 5$ is a prime and that $F$ contains a primitive $m$th root of unity. Let $a\in K\smallsetminus F$. Then $K[t;\sigma]/K[t;\sigma](t^{m}-a)$ is a division algebra if and only if
@@
Let $K/F$ be a cyclic Galois extension of $F$ of degree $n$ and let $\sigma$ be a generator of $\mathrm{Gal}(K/F)$. Suppose either that $m=2$ or $3$, or that $m\geq 5$ is a prime and $F$ contains a primitive $m$th root of unity.
@@
(ii) If $F$ is a finite field and $N_{K/F}(a)\not\in(F^{\times})^{m}$, then $K[t;\sigma]/K[t;\sigma](t^{m}-a)$ is a division algebra.
@@
For $n=m$, the unital central nonassociative algebra $K[t;\sigma]/K[t;\sigma](t^{m}-a)$ is called a nonassociative cyclic algebra of degree $n$ and denoted by $(K/F,\sigma,a)$ (cf. [Ste2014] who studies the opposite algebras, however). When $a\in F^{\times}$, the associative algebra $(K/F,\sigma,a)$ is a classical associative central simple cyclic algebra of degree $m$ [Jac1996, p. 19]. For $n=m$ we also have a stronger result than Corollary 2.6:
@@ 3. Different generators of $\mathrm{Gal}(K/F)$ yield nonisomorphic Sandler algebras
Let $F$ be an arbitrary field such that $F$ admits a cyclic Galois extension $K$ of degree $n>2$ and let $\sigma$ be a generator of the Galois group of $K/F$. We will now generalize Theorem [NevPu2024, Theorem 5.1] which covers the case $n=m$ that occurs when studying nonassociative cyclic algebras $(K/F,\sigma,a)$ to the case that $n\geq m-1$, starting with some general observations.
@@
Let $\sigma_{1}$ and $\sigma_{2}$ be any two distinct generators of $\mathrm{Gal}(K/F)$, i.e. There exists $1<j<n$, with $(j,n)=1$, such that $\sigma_{1}=\sigma_{2}^{j}$. Let $f\in K[t;\sigma_{1}]$ and $g\in K[t;\sigma_{2}]$ be both not right-invariant of degree $m$, and suppose that $n\geq m-1$. Let
@@
for $i\in\{0,\ldots,m-1\}$. Since $\sigma_{1}=\sigma_{2}^{j}$ and $\sigma_{i}$ has order $n\geq m-1$ it follows that $H(t)=kt^{j}$ for some $k\in K$. Since $H$ is an isomorphism, we know $k\in K^{\times}$. ∎
@@
Let $\sigma_{1}$ and $\sigma_{2}$ be any two distinct generators of $\mathrm{Gal}(K/F)$. Let $t^{m}-a_{1}\in K[t;\sigma_{1}]$ and $t^{m}-a_{2}\in K[t;\sigma_{2}]$ be not right-invariant, and suppose that $n\geq m-1$. Then
@@
Since $\sigma_{1},\sigma_{2}$ both generate $\mathrm{Gal}(K/F)$, there is some $1<j<n$, with $(j,n)=1$, such that $\sigma_{1}=\sigma_{2}^{j}$. Let $\ell$ be the least positive integer such that $\ell j>m$; since $\sigma_{1}\neq\sigma_{2}$, $\ell<m$. Suppose that $H:K[t;\sigma_{1}]/K[t;\sigma_{1}](t^{m}-a_{1})\rightarrow K[t;\sigma_{2}]/K[t;\sigma_{2}](t^{m}-a_{2})$ is an $F$-algebra isomorphism. Then $H|_{K}=\tau\in{\rm Gal}(K/F)$ and $H(t)=kt^{j}$ for some $k\in K^{\times}$ by Lemma 3.2. Since the Galois group is cyclic, and thus abelian, $\tau$ commutes with $\sigma_{i}$ on $K$, for $i\in\{1,2\}$. For any $s\geq 1$, put $\gamma_{s}(k)=k\sigma_{2}^{j}(k)\sigma_{2}^{2j}(k)\cdots\sigma_{2}^{(s-1)j}(k)$. Then one can show by induction that
$H(t^{s})=\begin{cases}\gamma_{s}(k)t^{sj}&\text{if $1\leq s<\ell$;}\\ \gamma_{\ell}(k)\sigma^{\ell j-m}(a_{2})t^{\ell j-m}&\text{if $s=\ell$.}\end{cases}$
@@
$t\cdot t^{\ell}=t^{\ell}\cdot t=\begin{cases}t^{\ell+1}&\text{if $\ell+1<m$};\\ a_{1}&\text{if $\ell+1=m$.}\end{cases}$
@@
$H(t^{\ell})H(t)=\gamma_{\ell}(k)\sigma^{\ell j-m}(a_{2})t^{\ell j-m}\cdot kt^{j}$
@@
$H(t^{\ell})H(t)=\begin{cases}\gamma_{\ell+1}(k)\sigma^{\ell j-m}(a_{2})t^{\ell j+j-m}&\text{if $(\ell+1)j<2m$;}\\ \gamma_{\ell+1}(k)\sigma^{\ell j-m}(a_{2})\sigma^{{\ell j+j-2m}}(a_{2})t^{\ell j+j-2m}&\text{if $2m\leq(\ell+1)j$}\end{cases}$
@@
$H(t)H(t^{\ell})=kt^{j}\cdot\gamma_{\ell}(k)\sigma_{2}^{\ell j-m}(a_{2})t^{\ell j-m},$
@@
$H(t)H(t^{\ell})=\begin{cases}\gamma_{\ell+1}(k)\sigma_{2}^{\ell j-m+j}(a_{2})t^{\ell j+j-m}&\text{if $(\ell+1)j<2m$;}\\ \gamma_{\ell+1}(k)\sigma_{2}^{\ell j-m+j}(a_{2})\sigma_{2}^{\ell j-2m+j}(a_{2})t^{\ell j+j-2m}&\text{if $2m\leq(\ell+1)j$}.\end{cases}$
@@
Let $n\geq m-1$. Let $\varphi$ be the Euler function. Then there are $\varphi(n)$ different classes of proper Petit algebras of type $K[t;\sigma_{i}]/K[t;\sigma_{i}](t^{m}-a)$; one class for each choice of generator $\sigma_{i}$ of $\mathrm{Gal}(K/F)$. Algebras in two different classes are not isomorphic.
The assumption that $n\geq m-1$ in Lemma 3.2 is indeed needed in the proof for (3.3). If $n<m-1$ we obtain a weaker version:
@@
Let $F$ be an arbitrary field and let $n\geq 3$ be the degree of a cyclic Galois extension $K/F$. Let $\sigma_{1}$ and $\sigma_{2}$ be any two distinct generators of the Galois group $\mathrm{Gal}(K/F)$, i.e. there is some $1<j<n$, with $(j,n)=1$, such that $\sigma_{1}=\sigma_{2}^{j}$. Let $f\in K[t;\sigma_{1}]$ and $g\in K[t;\sigma_{2}]$ be both not right-invariant of degree $m$, $n<m-1$. Let